can you replace the ques. marks with math symbol (-/*) and make the question correct (9?1)?(7?6)?(8?4)=3
Answer by Edwin McCravy(20055)

Make them all minuses:

(9-1)-(7-6)-(8-4) = 8-1-4 = 3

Edwin